40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Mar 4 13:37:14 2021, aaron, DailyProgress, Electronics, PSOMA S mA->MHz, ITC 502 current noise mA_to_MHz.pdf002F0263-0615-4566-9192-F57A14C8B543.jpegITC502_ASD.pdf
    Reply  Sun Mar 7 19:32:32 2021, rana, DailyProgress, Electronics, PSOMA S mA->MHz, ITC 502 current noise 
    Reply  Wed Mar 10 16:03:19 2021, aaron, DailyProgress, Electronics, current noise 
       Reply  Thu Mar 11 13:56:47 2021, aaron, DailyProgress, Electronics, current noise wire_shelf_5.jpg
Message ID: 2675     Entry time: Sun Mar 7 19:32:32 2021     In reply to: 2673
Author: rana 
Type: DailyProgress 
Category: Electronics 
Subject: PSOMA S mA->MHz, ITC 502 current noise 

something weird in the current noise plot - left axis log, right axis linear

also, something is up with the overall scale. With 5 nV of noise on a 20 Ohm resistor, the noise floor above 10 Hz should be ~250 pA/rHz. I recammend plotting the noise floor of the measurement apparatus with a 20 or 50 Ohm input term on the 560 (and grab the Busby Low Noise box from the 40)

ELOG V3.1.3-